# figurec (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v1

@@ 1 Introduction
Video generation has made tremendous progress in recent years. Results from Sora [7], OpenAI’s recently released text-to-video generation model, have shown that generating a high-quality video as long as one minute is possible. Following the scaling curve, video models will most likely continue to improve in many aspects. However, one essential functionality is still missing for these video models to be useful for many downstream applications – the ability to generate the same dynamic scene from an arbitrary camera perspective.
In this paper, we aim to tackle the problem of dynamic novel view synthesis (DVS) – given a video of a dynamic scene, we aim to generate a video of the same scene from another viewpoint. Once we develop a solution for this problem, we can leverage it for several impactful use cases, such as generating novel view of a live street scenario based on cameras mounted on an autonomous vehicle; seeing a cluttered environment from a different viewpoint while a robot is performing dexterous manipulation; and reliving a video recorded in the past from different viewing angles to make it more immersive.
@@
Free-viewpoint synthesis from a single video requires prior knowledge because it is highly under-constrained. Modern video generative models, such as Stable Video Diffusion [5], have learned rich priors for real-world dynamics, 3D geometry, and camera motions, as they are trained on hundreds of millions of video clips from the Internet. In this work, we propose an approach to capitalize on these rich representations for the task of DVS. We curate pairs of videos of dynamic scenes from simulation as training data, and apply them to steer a pretrained video generative model by means of finetuning.
@@ 2 Related Work
We note that at least one concurrent work also tackles dynamic view synthesis: in Exo2Ego [37], authors propose a framework that translates third-person (exocentric) to first-person (egocentric) videos, incorporating priors for hand-object interactions and focusing primarily on those scenarios.
Figure 2: Method. Our model, GCD, embodies an end-to-end video translation pipeline that maps an input video from any viewpoint into an output video from any other perspective, with the objective of respecting all objects and dynamics occurring within the observed dynamic scene, and faithfully reconstructing the corresponding visual details from this novel viewpoint. The relative camera extrinsics matrix $\Delta E$ guides the relationship between the two camera poses.
@@ 3 Approach
First, we formally introduce the task of monocular dynamic novel view synthesis from unconstrained video input. Let $\boldsymbol{x}\in\mathbb{R}^{T\times H\times W\times 3}$ be RGB frames captured from a single camera perspective, that encode the visual observation of a dynamic scene of interest. We denote its associated camera extrinsics matrix as $\mathcal{E}_{src}\in\mathbb{R}^{T\times 3\times 4}$, and define $\mathcal{E}_{dst}\in\mathbb{R}^{T\times 3\times 4}$ to be the desired target camera extrinsics matrix over time. Our model $f$ is then tasked with predicting a video $\boldsymbol{y}\in\mathbb{R}^{T\times H\times W\times 3}$, that plausibly depicts the same dynamic scene from the specified new viewpoint. For simplicity, and without loss of generality, we assume that (1) the output video is temporally synchronized with the input, and (2) the camera intrinsics matrix $\mathcal{K}$ stays constant over time as well as across pose changes; notably, the virtual camera for $\boldsymbol{y}$ assumes the same focal length as the actual camera does for $\boldsymbol{x}$.
@@ 3.1 Camera viewpoint control
Given a single RGB video $\boldsymbol{x}$ of a dynamic scene, our goal is to synthesize another video $\boldsymbol{y}$ of the scene from a different viewpoint. Since large-scale video diffusion models have been trained on hyper-scale data, their support of the natural video distribution most likely covers a wide range of realistic scenes and viewpoints. To this end, given a dataset of paired videos and their relative camera extrinsics $\Delta\mathcal{E}=\{\mathcal{E}_{src,t}^{-1}\cdot\mathcal{E}_{dst,t}\}_{t=1}^{T}\in\mathbb{R}^{T\times 3\times 4}$ over time, we teach a latent diffusion model $f$ to learn controls over camera parameters within any video:
@@
Specifically, we modify Stable Video Diffusion (SVD) to accept a new form of micro-conditioning, a term coined in [5], which is designed for the purpose of communicating low-dimensional metadata (such as the desired frame rate of the output video, and the amount of optical flow) to the network. We decompose $\Delta\mathcal{E}$ into a series of camera rotation matrices $R_{t}\in\mathcal{SO}(3)$ and translation matrices $T_{t}\in\mathbb{R}^{3}$ over time, project this information through an MLP $m$, and add the resulting embedding to the feature vectors at various convolutional layers placed throughout the network, similarly to the concurrent work SV3D [64]. The diffusion timestep, FPS, and motion strength are also passed to the network this way. To preserve the existing priors of SVD as much as possible, we initialize the network weights based on the publicly available image-to-video model checkpoint. The new embedder $m$ that processes $\{(R_{t},T_{t})\}$ is randomly initialized with default parameters. After training the network end-to-end to tackle this new task, the resulting model is capable of imagining unseen videos from any chosen perspective, as illustrated in Figure 2.
@@ 3.2 Video conditioning
To accurately perform 4D dynamic scene completion, both low-level perception (to analyze the visible shapes, appearance, etc.) and high-level understanding (to infer the occluded regions, based on world knowledge as well as other observed frames) of the input video is required. We adopt the same hybrid conditioning mechanism as SVD [5], where the visual signal is processed in two ways. In case of image-to-video, the first stream calculates the CLIP [47] embedding $c(\boldsymbol{x}_{0})$ of the incoming image to condition the U-Net $\epsilon$ via cross-attention, and the second stream channel-concatenates the VAE-encoded image $\boldsymbol{x}_{0}$ with all frames of the video sample $\hat{\boldsymbol{y}}$ that is being denoised by the diffusion model.
@@
In architectural terms, the output video sample $\hat{\boldsymbol{y}}$ has contemporaneous input frames from $\boldsymbol{x}$ attached to it for every video timestamp $t$, such that at diffusion noise timestep $u$:
$\hat{\boldsymbol{y}}_{u-1}=\epsilon\left(\hat{\boldsymbol{y}}_{u}\mathbin{\|}\boldsymbol{x},\Delta\mathcal{E}\right),$ (2)
In other words, the U-Net $\epsilon$ accepts input feature maps of dimensionality $2D\times T\times\frac{H}{F}\times\frac{W}{F}$, where $D$ and $F$ are the VAE embedding size and downsampling factor respectively, and produces output feature maps of dimensionality ${D\times T\times\frac{H}{F}\times\frac{W}{F}}$ that represent a less noisy sample.
Note that the SVD architecture consists of a factorized 3D U-Net that interleaves spatial and temporal blocks, which establish correspondences between features across locations (per frame), and across time (per spatial position) respectively. Spatiotemporal attention can consequently take place between all pairs of input and output frames, as well as any pair of regions within both videos. Moreover, there are now $T$ different CLIP embeddings $\{c(\boldsymbol{x}_{t})\}$ that appropriately condition the U-Net layers at each matching frame.
@@ 4 Datasets
While the availability of multi-view video data has been growing [13, 62, 52, 19, 61, 48, 87, 20], it is still relatively sparse compared to conventional image or video datasets. In order to train and evaluate our model, we require a decent amount of multi-view RGB videos from highly cluttered dynamic scenes. To this end, we contribute two high-quality synthetic datasets, shown in Figures 34, and 5 and briefly described below.
@@ 4.3 Task details
When training for the task of dynamic view synthesis on Kubric-4D, pairs of input and output poses are randomly sampled within certain spherical coordinate bounds (both in absolute terms and relative to each other), with the extra condition that they are looking at the center of the 3D scene. Therefore, at inference time, there are three effective degrees of freedom with regard to camera control, similar to Zero-1-to-3[34].
@@ 5 Choice of camera trajectory
Our formulation of the dynamic view synthesis task in Section 3 is quite general, so it is worth thinking about which specific instantiations of this conceptual framework would be most effective in practice. Given arbitrary video inputs, our goal is to devise a structured protocol for choosing relative camera trajectories that both maximize the exploitation of knowledge contained within the pretrained SVD representation, as well as enable a detailed understanding of the dynamic scene observed at inference time to the fullest extent possible. Specifically, we wish to synthesize views that reach as far as the opposite end of the scene, for example, by rotating the azimuth angle up to 180°. This is considerably more dramatic than what the state of the art in dynamic view synthesis is typically capable of [17, 10, 31, 74], and allows us to reveal large, formerly unseen portions of the surroundings.
@@
To resolve this concern, we translate it into three questions: (1) where should the output pose start; (2) how fast should it be taught to move in-between subsequent frames; and (3) how much does finetuning, i.e. borrowing priors from SVD help (or hurt) in each case, versus training an identical network from scratch? We investigate this by running comparative studies on both the Kubric-4D and ParallelDomain-4D datasets. For each tested scene, we pick a source pose $\mathcal{E}_{{src},0}$ and a destination pose $\mathcal{E}_{{dst},T}$. For our experiments, we assume a static input camera, i.e., $\mathcal{E}_{src}=\{\mathcal{E}_{{src},0}\}_{t=1}^{T}$.
The results are shown in Tables 12, and 3, and Figures 3 and 4. Here, gradual means that the virtual camera pose corresponding to the output video linearly interpolates between $\mathcal{E}_{{src},0}$ and $\mathcal{E}_{{dst},T}$ from start to end (in spherical coordinates), whereas direct implies that the generated video directly adheres precisely to $\mathcal{E}_{{dst},T}$ at every frame without interpolation. For Kubric-4D, max 90° limits the relative horizontal (i.e. azimuth) angle variation between input and output to ±$\pm 90\degree$ at training time, whereas max 180° effectively allows for synthesizing any 360°-surround viewpoint of the dynamic scene.
From this ablation study, we observe with Kubric-4D that (1) it is preferable to gradually interpolate from source to destination pose than to immediately jump there (+1.17 dB average PSNR improvement between direct and gradual); (2) there exists a trade-off between the range of camera transformations the model should be trained for, and how extreme of a rotation one wishes to be able to achieve at most (+0.55 dB between max 180° and max 90°); and (3) it is preferable to start from the SVD checkpoint that had been trained on large-scale video rather than to train from random initialization, though not by a particularly huge margin (+1.34 dB between scratch and finetuned).
@@ 6.1 Implementation details
We adopt the SVD variant that predicts $T=14$ frames, but due to computational constraints, we downscale the input and output resolution to $W\times H=384\times 256$. This allows us to scale the batch size up to 56 when training with Kubric-4D on 7x A100 GPUs with 80 GB VRAM each. We finetune all models for 10k iterations using the Adam optimizer, which takes roughly 3 days. On ParallelDomain-4D, we instead finetune models for 13k iterations with an effective batch size of 24 through a gradient accumulation factor of 4 on 7x A6000 GPUs with 48 GB VRAM each, which also takes roughly 3 days.
@@
We generate conditional samples from the resulting diffusion model by running the EDM sampler for 25 steps [27]. SVD originally employs classifier-free guidance at test time with a guidance strength $w$ that linearly increases as a function of the video frame index (not the diffusion timestep) from start to end within the range $[1,2.5]$ by default [5], but we found better performance by adjusting this range to $[1,1.5]$ instead. Producing one output video takes roughly 10 seconds.
@@
Following related work in novel view synthesis [38, 28, 50, 34, 31, 67], for predictions in RGB space, we evaluate PSNR, SSIM, and LPIPS scores and average the results across both video frames and test examples. For semantic category predictions, following conventions in semantic segmentation [12, 77, 57, 86], we first calculate the average Intersection over Union (IoU) per category over the whole ParallelDomain test set, and then report the mean IoU (mIoU) across categories.
@@ 6.3 Results
We report quantitative results in Tables 45, and 6, and show qualitative results in Figures 5 and 4. On both datasets, our model outperforms baseline methods by a large margin. Per-scene optimization methods (e.g., HexPlane) fail to reconstruct the 4D scene representation from a single input view, and thus the rendered videos from novel viewpoints have severe artifacts. Vanilla SVD is able to generate smooth videos but fails to follow the desired camera trajectories, and does not incorporate content from later frames. ZeroNVS can synthesize plausible individual frames from desired viewpoints, but the resulting videos are not temporally coherent and do not respect the scene dynamics.
In contrast, our model mostly generates plausible videos that accurately depict the complex scene geometry and motion under the desired novel viewpoint transformations. We remark that the results are not perfect, as the correspondence of objects between the input and generated output videos is not always very clear, and some rigid objects tend to erroneously deform. However, the nature of the task is extremely challenging, and we expect the potential visual quality and consistency of our framework to only improve over time, e.g. when combined with better pretrained representations.

### v2

@@ 1 Introduction
Video generation has made tremendous progress in recent years. Results from Sora [7], OpenAI’s recently released text-to-video generation model, have shown that generating a high-quality video as long as one minute is possible. Following the scaling curve, video models will most likely continue to improve in many aspects. However, one essential capability is still missing from these video models to be useful for many downstream applications – the ability to generate the same dynamic scene from an arbitrary camera perspective based on an existing video.
In this paper, we aim to tackle the problem of dynamic novel view synthesis (DVS) – given a video of a dynamic scene, we aim to generate a video of the same scene from another viewpoint. Once we develop a solution for this problem, we can leverage it for several impactful use cases, such as generating novel views of a live street scenario based on cameras mounted on an autonomous vehicle; seeing a cluttered environment from a different viewpoint while a robot is performing dexterous manipulations; enabling geometrically consistent video passthrough for mixed reality [81]; and immersively reliving videos recorded in the past from different viewing angles.
@@
Free-viewpoint synthesis from a single video requires prior knowledge because it is highly under-constrained. Modern video generative models, such as Stable Video Diffusion [5], have learned rich priors for real-world dynamics, 3D geometry, and camera motions, as they are trained on hundreds of millions of video clips from the Internet. In this work, we propose an approach to capitalize on these rich representations for the task of DVS. We curate pairs of videos of dynamic scenes from simulation as training data, and apply them to steer a pretrained video generative model towards the desired behavior by means of finetuning.
@@ 2 Related Work
We note that at least one concurrent work also tackles dynamic view synthesis: in Exo2Ego [39], authors propose a framework that translates third-person (exocentric) to first-person (egocentric) videos on a per-frame basis, incorporating priors for hand-object interactions and focusing primarily on those scenarios.
Figure 2: Method. Our model, GCD, is an end-to-end video translation pipeline that maps an input video from any viewpoint into an output video from any other perspective, with the objective of respecting all objects and dynamics occurring within the observed dynamic scene, and faithfully reconstructing the corresponding visual details from this novel viewpoint. The relative camera extrinsics matrix $\Delta\mathcal{E}$ guides the relationship between the two camera poses.
@@ 3 Approach
First, we formally introduce the task of monocular dynamic novel view synthesis from unconstrained video input. Let $\boldsymbol{x}\in\mathbb{R}^{T\times H\times W\times 3}$ be RGB frames captured from a single camera perspective, that encode the visual observation of a dynamic scene of interest. We denote its associated camera extrinsics matrix as $\mathcal{E}_{src}\in\mathbb{R}^{T\times 4\times 4}$, and define $\mathcal{E}_{dst}\in\mathbb{R}^{T\times 4\times 4}$ to be the desired target camera extrinsics matrix over time. Our model $f$ is then tasked with predicting a video $\boldsymbol{y}\in\mathbb{R}^{T\times H\times W\times 3}$, that plausibly depicts the same dynamic scene from the specified new viewpoint. For simplicity, and without loss of generality, we assume that (1) the output video is temporally synchronized with the input, and (2) the camera intrinsics matrix $\mathcal{K}\in\mathcal{R}^{3\times 3}$ stays constant over time as well as across pose changes; notably, the virtual camera for $\boldsymbol{y}$ assumes the same focal length as the actual camera does for $\boldsymbol{x}$.
@@ 3.1 Camera viewpoint control
Given a single RGB video $\boldsymbol{x}$ of a dynamic scene, our goal is to synthesize another video $\boldsymbol{y}$ of the scene from a different viewpoint. Since large-scale video diffusion models have been trained on hyper-scale data, their support of the natural video distribution most likely covers a wide range of realistic scenes and viewpoints. To this end, given a dataset of paired videos and their relative camera extrinsics $\Delta\mathcal{E}=\{\mathcal{E}_{src,t}^{-1}\cdot\mathcal{E}_{dst,t}\}_{t=0}^{T-1}\in\mathbb{R}^{T\times 4\times 4}$ over time, we teach a latent diffusion model $f$ to learn controls over camera parameters within any video $\boldsymbol{x}$:
@@
Specifically, we modify Stable Video Diffusion (SVD) to accept a new form of micro-conditioning, a term coined in [5], which is designed for the purpose of communicating low-dimensional metadata (such as the desired frame rate of the output video, and the amount of optical flow) to the network. We decompose $\Delta\mathcal{E}_{t}\in\mathrm{SE}(3)$ into a series of camera rotation matrices $R_{t}\in\mathrm{SO}(3)$ and translation matrices $T_{t}\in\mathbb{R}^{3}$ over time, project this flattened information through an MLP $m$, and add the resulting embedding to the feature vectors at various convolutional layers placed throughout the network, similarly to the concurrent work SV3D [68]. The diffusion timestep, FPS, and motion strength are also passed to the network this way. To preserve the existing priors of SVD as much as possible, we initialize the network weights based on the publicly available image-to-video model checkpoint. The new embedder $m$ that processes $\{(R_{t},T_{t})\}$ is randomly initialized with default parameters. After training the network end-to-end to tackle this new task, the resulting model is capable of imagining unseen videos from any chosen perspective, as illustrated in Figure 2 (high-level) and Figure 10 (detail).
@@ 3.2 Video conditioning
To accurately perform dynamic view synthesis, both low-level perception (to analyze the visible geometry, shapes, appearance, etc.) and high-level understanding (to infer the occluded regions, based on world knowledge as well as other observed frames) of the input video is required. We adopt the same hybrid conditioning mechanism as SVD [5], where the visual signal is processed in two ways. In case of image-to-video, the first stream calculates the CLIP [49] embedding $c(\boldsymbol{x}_{0})$ of the incoming image to condition the U-Net $\epsilon$ via cross-attention, and the second stream channel-concatenates the VAE-encoded image $\boldsymbol{x}_{0}$ with all frames of the video sample $\hat{\boldsymbol{y}}$ that are being denoised.
@@
In architectural terms, the output sample $\hat{\boldsymbol{y}}$ has contemporaneous input frames from $\boldsymbol{x}$ attached to it for every video timestamp $t$, such that at diffusion noise timestep $u$ during inference:
$\hat{\boldsymbol{y}}_{u-1}=w\epsilon\left(\hat{\boldsymbol{y}}_{u}\mathbin{\|}\boldsymbol{x},\Delta\mathcal{E}\right)-(w-1)\epsilon\left(\hat{\boldsymbol{y}}_{u}\right),$ (2)
where $w\in[1,\infty)$ is the guidance strength for classifier-free guidance [25]. The U-Net $\epsilon$ accepts input feature maps of dimensionality $2D\times T\times\frac{H}{F}\times\frac{W}{F}$, where $D$ and $F$ are the VAE embedding size and downsampling factor respectively, and produces output feature maps of dimensionality ${D\times T\times\frac{H}{F}\times\frac{W}{F}}$ that represent a less noisy sample.
Note that the SVD architecture consists of a factorized 3D U-Net that interleaves convolutional, spatial, and temporal blocks, of which the latter two establish correspondences between features across locations (per frame), and across time (per spatial position) respectively. Spatiotemporal attention can consequently take place between all pairs of input and output frames, as well as any pair of regions within both videos. Moreover, there are now $T$ different CLIP embeddings $\{c(\boldsymbol{x}_{t})\}$ that appropriately condition the U-Net layers at each matching frame.
@@ 4 Datasets
While the availability of multi-view video data has been growing [13, 66, 56, 19, 65, 50, 92, 20, 30, 51], it is still relatively sparse compared to conventional image or video datasets. In order to train and evaluate our model, we require a decent amount of multi-view RGB videos from highly cluttered dynamic scenes. To this end, we contribute two high-quality synthetic datasets, shown in Figures 34, and 5 and briefly describe them below.
@@ 4.3 Task details
In our experiments, without loss of generality, we assume a static input camera pose $\mathcal{P}_{{src}}$, [^1] This can always be achieved by defining the reference coordinate system to move along with the recording camera. and pick a target destination pose $\mathcal{P}_{{dst}}$ that we want the output camera to reach at or before the end of the generated video. In general, $\mathcal{P}$ is a pose description that can be defined in many ways, for example a set of spherical coordinates that represent the camera position and look-at location, but (1) must allow for convex interpolation (i.e. $\alpha\mathcal{P}_{1}+(1-\alpha)\mathcal{P}_{2}$ with $\alpha\in[0,1]$ is valid), and (2) is connected to a valid 6-DoF rigid body transformation $\mathcal{E}\in\mathrm{SE}(3)$ through the function $g$, i.e. $\mathcal{E}=g(\mathcal{P})$.
When training for the task of dynamic view synthesis on Kubric-4D, pairs of input and output poses are randomly sampled within certain spherical coordinate bounds (both in absolute terms and relative to each other), with the extra condition that they are looking at the center of the 3D scene. Therefore, at inference time, there are three effective degrees of freedom with regard to camera control, i.e. $\mathcal{P}\in\mathbb{R}^{3}$.
@@ 5 Choice of camera trajectory
Our formulation of the dynamic view synthesis task in Section 3 is quite general, so it is worth thinking about which specific instantiations of this conceptual framework would be most effective in practice. Given arbitrary video inputs, our goal is to devise a structured protocol for choosing relative camera trajectories that both maximize the exploitation of knowledge contained within the pretrained SVD representation, as well as enable a detailed understanding of the dynamic scene observed at inference time to the fullest extent possible. Specifically, we wish to synthesize views that reach as far as the opposite end of the scene, for example, by orbiting the azimuth angle $\phi$ up to 180°. This is considerably more dramatic than what the state of the art in dynamic view synthesis is typically capable of [17, 10, 33, 78], and allows us to reveal large, formerly unseen portions of the surroundings.
@@
To resolve this concern, we translate it into three questions: (1) where should the output pose start; (2) how fast should it be taught to move in-between subsequent frames; and (3) how much does finetuning, i.e. borrowing priors from SVD help (or hurt) in each case, versus training an identical network from scratch? We investigate this by running comparative studies on both the Kubric-4D and ParallelDomain-4D datasets. For each tested scene, we fix a source pose $\mathcal{P}_{{src}}$ and a destination pose $\mathcal{P}_{{dst}}$, following definitions in Section 4.3. Using $\mathcal{E}_{src,t}=g(\mathcal{P}_{src})$ and $\alpha=\frac{t}{T-1}\in[0,1]$, we define gradual and direct trajectories as follows:
$\mathcal{E}_{dst,t}=\begin{cases}g\left(\alpha\mathcal{P}_{dst}+(1-\alpha)\mathcal{P}_{src}\right),&\forall t,\hskip 11.38092pt\text{if gradual}\\ g\left(\mathcal{P}_{{dst}}\right),&\forall t,\hskip 11.38092pt\text{if direct}\end{cases}$ (3)
In other words, gradual means that the virtual camera pose corresponding to the output video linearly interpolates (in an intermediate description space, for example spherical coordinates) between $\mathcal{P}_{{src}}$ and $\mathcal{P}_{{dst}}$ from start to end, whereas direct implies that the generated video directly adheres precisely to $\mathcal{P}_{{dst}}$ at every frame without interpolation. For Kubric-4D, max 90° limits the relative horizontal (i.e. azimuth) angle variation between input and output to ±$\pm 90\degree$ at training time ($|\Delta\phi|\leq 90\degree$), whereas max 180° effectively allows for synthesizing any 360°-surround viewpoint of the dynamic scene.
The results are shown in Tables 12, and 3, and Figures 3 and 4. From this ablation study, we observe with Kubric-4D that: (1) it is preferable to gradually interpolate from source to destination pose than to immediately jump there (+1.17 dB average PSNR improvement between direct and gradual); (2) there exists a trade-off between the range of camera transformations the model should be trained for, and how extreme of a rotation one wishes to be able to achieve at most (+0.55 dB between max 180° and max 90°); and (3) it is preferable to start from the SVD checkpoint that had been trained on large-scale video rather than to train from random initialization, though not by a particularly huge margin (+1.34 dB between scratch and finetuned).
@@ 6.1 Implementation details
We adopt the SVD variant that predicts $T=14$ frames, but due to computational constraints, we downscale the input and output resolution to $W\times H=384\times 256$. This allows us to scale the batch size up to 56 when training with Kubric-4D on 7x A100 GPUs with 80 GB VRAM each. We finetune all models for 10k iterations using the Adam optimizer, which takes roughly 3 days. On ParallelDomain-4D, we instead finetune models for 13k iterations with an effective batch size of 24 through a gradient accumulation factor of 4 on 3x A6000 GPUs with 48 GB VRAM each, which also takes roughly 3 days. The network $\epsilon$ does not predict noise directly, instead adopting v-parameterization for preconditioning [53].
@@
We generate conditional samples from the resulting diffusion model by running the EDM sampler [28] for 25 steps. SVD originally employs classifier-free guidance [25] at test time with a guidance strength $w$ that linearly increases as a function of the video frame index (not the diffusion timestep) from start to end within the range $[1,2.5]$ by default [5], but we found better performance by adjusting this range to $[1,1.5]$ instead. Producing one output video takes roughly 10 seconds.
@@
Following related work in novel view synthesis [40, 29, 54, 36, 33, 71], for predictions in RGB space, we evaluate PSNR, SSIM, and LPIPS scores and average the results across both video frames and test examples. For semantic category predictions, following conventions in semantic segmentation [12, 82, 61, 91], we first calculate the average Intersection over Union (IoU) per category over the whole ParallelDomain test set, and then report the mean IoU (mIoU) across the 10 most common categories.
@@ 6.3 Results
We report quantitative results in Tables 45, and 6, and show qualitative results in Figures 5 and 4. On both datasets, our model outperforms baseline methods by a large margin. Per-scene optimization methods (e.g., HexPlane) fail to reconstruct the 4D scene representation from a single input view, and thus the rendered videos from novel viewpoints have severe artifacts. Vanilla SVD is able to generate smooth videos but fails to follow the desired camera trajectories, and does not incorporate content from later frames. ZeroNVS can synthesize plausible individual frames from specified viewpoints, but the resulting videos are not temporally coherent and do not respect the scene dynamics.
In contrast, our model mostly generates plausible videos that accurately depict the complex scene geometry and motion under the desired novel viewpoint transformations. We remark that the results are not perfect, as the correspondence of objects between the input and generated output videos is not always very clear, and some rigid objects tend to erroneously deform. However, the nature of the task is extremely challenging, and we expect the potential visual quality and consistency of our framework to only improve over time, e.g. when combined with better pretrained representations, more compute, and more data.